# Measuring the Dunkelflaute: How (not) to analyze variable renewable energy shortage

### v2

@@ Abstract
journal: arXiv
@@ 1 Introduction
Another approach is to consider not only VRE generation potentials, but also how these relate to electric load profiles. This can be done by analyzing residual load, sometimes also referred to as net load, which represents the net balance of electricity demand and VRE supply. Again, there is no consensus in the literature on how to term such shortage events (Figure 1), which, for example, have been referred to as “energy supply droughts” [19], “energy deficits” [22], or “energy shortfall events” [23]. For clarity, we propose the following terminology: “variable renewable energy droughts” for periods with low resource availability of a single or multiple VRE technologies, “variable renewable energy anomaly” for cumulative deviations of VRE resource availability from some mean reference, and “ positive residual load (PRL) events” for periods where variable renewable supply falls short of electric load, and “variable renewable energy shortage” as a generic term that captures these concepts (Figure 1). Note that VRE drought and PRL events refer to concrete periods and do not necessarily occur at the same time. While both concepts characterize variable renewable energy shortage, PRL events are additionally affected by the timing of coinciding high-demand periods. In contrast, VRE anomalies are measured over user-specific time horizons.
@@
Different bodies of literature contribute to the analysis of VRE shortage events, primarily related to meteorology or energy research. One literature strand focuses on wind energy. Notably, wind droughts in the United Kingdom are well-studied. Cannon et al. [24] explore extreme onshore wind power events in the UK based on reanalysis data covering 33 years. Thresholds correspond to percentiles of the cumulative frequency distribution of investigated data. Potisomporn & Vogel [25] study the spatial and temporal correlation of low-wind-power events for offshore wind locations in the United Kingdom given a historical and projected distribution of offshore wind farms. Potisomporn et al. [26] contrast low-wind speed events derived from ERA5 reanalysis data to measured data from 205 onshore and offshore wind locations in the United Kingdom. The same authors [17] compare different identification methods for low-wind events in the United Kingdom based on 72 years of ERA5 reanalysis data, deploying concepts for analyzing hydrological droughts [27, 28]. Abdelaziz et al. [29] assess offshore wind speed droughts in the United Kingdom. They identify consecutive days with daily average wind speed below a given threshold. There are related studies analyzing other regions. Leahy & McKeogh [30] and Patlakas et al. [31] study the persistence of low-wind speeds for different locations in Ireland and the North Sea based on historical records. They search for events with wind speeds constantly below a threshold and for annual minimal values of moving average wind speeds over pre-defined intervals, applying extreme value theory. Thresholds are based on a range of exogenously set wind speeds. Ohlendorf & Schill [32] investigate the duration and frequency of low onshore wind power events in Germany based on reanalysis data and exogenously set thresholds.
@@
While the first VRE drought analyses focused on wind power, research interest in droughts of variable renewable energy portfolios that also include solar photovoltaics (PV) is growing. Raynaud et al. [19] investigate droughts for wind, solar, and hydro, both separately and combined as a portfolio, for 12 European regions and 30 years of data. They employ relative regional thresholds as fractions of the mean daily availability factor. Rinaldi et al. [18] draw on Raynaud et al. [19] for identifying wind, solar, and wind-solar droughts in the Western U.S. for 39 years of data for different thresholds relative to mean daily generation. Gangopadhyay et al. [33] apply a similar approach to India, drawing on 5000 years of synthetic data. Kies et al. [34] compare seven VRE availability datasets for Europe. Using a composite VRE portfolio time series that combines wind and PV availability, they search for periods constantly below a range of exogenous thresholds. Kapica et al. [35] compare energy droughts for solar PV and wind in Europe, either separately or as a combined portfolio, for historical and future weather conditions retrieved from eight regional climate models. Their identification method again draws on [19] with thresholds relative to seasonal and regional averages. Hu et al. [36] explore the risk of simultaneously occurring VRE droughts for European countries for both single technologies and a VRE portfolio, and compare it to droughts arising across all of Europe assuming perfect electricity exchange. Breyer et al. [37] model decarbonization pathways for Europe until 2040 or 2050, using the weather year 2005. They screen their model outcomes for drought periods constantly below a range of thresholds relative to the maximum availability of wind and PV separately and combined for Germany and Europe. Kaspar et al. [38] analyze wind and PV separately and combined for Germany based on reanalysis data. For given averaging intervals, they scan for moving averages below exogenously set thresholds. Mockert et al. [39] also study VRE droughts in Germany in the context of prevailing weather regimes and lean on Kaspar et al. [38] for drought identification. Mayer et al. [40] analyze concurrent PV and wind droughts in Hungary by identifying consecutive time steps with availability from both technologies below exogenous thresholds. Ohba et al. [41] link wind and solar droughts, both separately and combined as a portfolio, to four typical weather patterns arising in eastern Japan using an exogenously set threshold and neural network learning.
@@
In another strand of the literature, VRE shortage events are analyzed in the context of residual load. In addition to their VRE drought analysis, Raynaud et al. [19] investigate the imbalance of demand and VRE supply in Europe. Van der Wiel et al. [23] test whether low wind and solar generation as well as high residual events are linked to large-scale weather regimes arising over Europe. Otero et al. [45] investigate periods with simultaneously low wind and solar generation and high demand across 27 European countries based on absolute thresholds. Otera et al. [42] use copula analysis to analyze VRE droughts based on a threshold relative to a ten percent fraction of its mean availability, and residual load events above the 90% percentile of its distribution in Europe based on ERA5 reanalysis. Allen & Otero [43] introduce standardized VRE droughts indices derived from meteorological indicators that incorporate both VRE generation and electricity demand. They showcase the proposed method for droughts in Europe. Ruhnau & Qvist [22] retrieve residual load time series from a model analysis of a least-cost decarbonized Germany power sector scenario and search for persistent positive residual load events. Dawkins & Rushby [8] characterize wind droughts and residual load events that qualify as droughts in the United Kingdom based on percentile thresholds. Francois et al. [46] explore positive residual load events (coined electricity deficits) for run-of-river hydropower and solar PV net of demand for Northern Italy. They deploy Complete-Time-Analysis, which is focused on return periods lasting from single hours up to several years. Based on the assumption of sufficient storage availability, Plain et al. [47] identify low-solar periods using 21 years of satellite data to size microgrids power systems in Africa based on the mismatch between solar production and demand. Bracken et al. [44] characterize VRE droughts across the U.S. concerning frequency, duration, magnitude, and seasonality given the existing energy infrastructure in 2020. They investigate single VRE technologies, VRE portfolios, and positive residual load events.
@@
Alternatively, Grochowicz et al. [48] detect power system stress events in Europe based on electricity shadow prices derived from a power sector model. They consider both large-scale weather patterns over 40 years of re-analysis data and power sector dynamics. For each year, they identify system-defining events that may last up to two weeks and account for a considerable share of total system costs. Subsequently, they classify the mid-troposphere dynamics into weather regimes and cluster near-surface meteorological conditions present during these events. System-defining events occur between November and February and last between two and 13 days. The analysis also takes into account the shadow prices of transmission capacity constraints and hourly storage energy constraints. Importantly, their findings are only applicable to the particular model setting and rely on a range of model-inherent parameter assumptions.
@@
We provide a structured overview of different approaches for identifying variable renewable energy droughts across multiple regions and technologies from the perspective of the energy system. Based on the literature review, we focus our analysis on threshold-based methods, including variable renewable energy droughts and positive residual load events. We discuss possible definitions of VRE droughts and identification methods, focusing on renewable availability time series. As variable renewable energy potentials vary across technologies, space, and time, we elaborate on how to enable comparability across these dimensions, which has not yet been considered in the literature. We further discuss how VRE drought analysis can be extended to also include electric load, i.e., how to evaluate extreme PRL. This allows for identifying compound low-VRE availability and high-demand events indicating the need for dispatchable capacity. Drawing on the proposed method for VRE availability time series, we discuss how to analyze PRL events and discuss essential points of criticism.
@@ 4.3 Methodological trade-offs and outlook
Such a trade-off also applies to the technological scope of VRE drought analysis. Aside from methodological assumptions inherent to the generation of variable renewable availability time series, e.g., such for reanalysis, insights from single-technology analysis are broadly applicable but less policy-relevant for future energy systems that rely on a mix of renewable technologies. Conversely, analyzing a VRE portfolio is more policy-oriented but involves energy modeling or exogenous assumptions on the composition of the investigated capacity mix. VRE capacity mixes determined by an energy modeling are optimal only in the context of the other optimized system components and flexibility options that are represented in the model (e.g., storage, electricity exchange, or direct-electrified applications from the heat, industry, or transport sector). Similarly, exogenous capacity mixes rely on user-specific assumptions. These aspects limit the validity of insights to investigated scenarios.
While we provide general guidance on the advantages and drawbacks of different VRE shortage identification methods, a thorough quantitative comparison would be desirable, ideally for a range of variable renewable technologies and electric load profiles for different world regions and weather years. Another key aspect that requires attention is the choice of the threshold parameter, particularly in distinguishing between severe VRE droughts and prolonged periods of below-average availability. This distinction could support the selection of relevant weather years to enable weather- and climate-resilient energy system analyses and well-informed decarbonization policies. Additionally, assessing the sensitivity of outcomes to different VRE availability datasets would be useful and could be valuable feedback to the meteorological research community that provides and continuously improves such datasets [50].

### v3

@@ Abstract
journal: Environmental Research: Energy
@@ 1 Introduction
Another approach is to consider not only VRE generation potentials, but also how these relate to electric load profiles. This can be done by analyzing residual load, sometimes also referred to as net load, which represents the net balance of electricity demand and VRE supply. Again, there is no consensus in the literature on how to term such shortage events (Figure 1), which, for example, have been referred to as “energy supply droughts” [19], “energy deficits” [22], or “energy shortfall events” [23].
Alternatively, power system model outcomes that use VRE availability as input can be used to identify extreme VRE conditions causing power system stress. Focusing on high electricity prices, corresponding periods have been termed as “system-defining” [24], “hydrometeorological compound events” [25], or “severe winter weather events” [26].
For clarity, we propose the following terminology: “variable renewable energy droughts” for periods with low resource availability of a single or multiple VRE technologies, “variable renewable energy anomaly” for cumulative deviations of VRE resource availability from a given mean reference, “ positive residual load (PRL) events” for periods where variable renewable supply falls short of electric load, and “electricity system stress events” for periods with high electricity prices signaling extreme renewable availability. The generic term “variable renewable energy shortage” captures these concepts (Figure 1). Note that VRE drought and PRL events refer to concrete periods and do not necessarily occur at the same time. While both concepts characterize variable renewable energy shortage, PRL events are additionally affected by the timing of coinciding high-demand periods. In contrast, VRE anomalies are measured over user-specific time horizons. Electricity system stress events may be driven by further technical limitations such as grid constraints, impacting their occurrence patterns.
@@
electricity system stress events
focusing on extreme electricity prices,
identified by electricity system models
system-defining events Grochowicz et al. [24]
hydrometeorological compound Su et al. [25]
events leading to extreme prices
severe winter weather events Akdemir et al. [26]
causing lower electricity generation
hydrometeorological extremes Wessel et al. [48]
causing extreme prices
@@
Different bodies of literature contribute to the analysis of VRE shortage events, primarily related to meteorology or energy research. One literature strand focuses on wind energy. Notably, wind droughts in the United Kingdom are well-studied. Cannon et al. [27] explore extreme onshore wind power events in the UK based on reanalysis data. Thresholds correspond to percentiles of the cumulative frequency distribution of investigated data. Potisomporn & Vogel [28] study the spatial and temporal correlation of low-wind-power events for offshore wind locations in the United Kingdom given a historical and projected distribution of offshore wind farms. Potisomporn et al. [29] contrast low-wind speed events derived from ERA5 reanalysis data to measured data from 205 onshore and offshore wind locations in the United Kingdom. The same authors in [17] compare different identification methods for low-wind events in the United Kingdom based on reanalysis data, deploying concepts for analyzing hydrological droughts [30, 31]. Abdelaziz et al. [32] assess offshore wind speed droughts in the United Kingdom. They identify consecutive days with daily average wind speed below a given threshold. There are related studies analyzing other regions. Leahy & McKeogh [33] and Patlakas et al. [34] study the persistence of low-wind speeds for different locations in Ireland and the North Sea based on historical records. They search for events with wind speeds constantly below a threshold and for annual minimal values of moving average wind speeds over pre-defined intervals. Thresholds are based on a range of exogenously set wind speeds. Ohlendorf & Schill [35] investigate the duration and frequency of low onshore wind power events in Germany based on reanalysis data and exogenously set thresholds.
@@
While the first VRE drought analyses focused on wind power, research interest in droughts of variable renewable energy portfolios that also include solar photovoltaics (PV) is growing. Raynaud et al. [19] investigate droughts for wind, solar, and hydro, both separately and combined as a portfolio, for 12 European regions. They employ relative regional thresholds as fractions of the mean daily availability factor. Rinaldi et al. [18] draw on Raynaud et al. [19] for identifying wind, solar, and wind-solar droughts in the Western U.S. for different thresholds relative to mean daily generation. Gangopadhyay et al. [36] apply a similar approach to India, drawing on 5000 years of synthetic data. Kies et al. [37] compare seven VRE availability datasets for Europe. Using a composite VRE portfolio time series that combines wind and PV availability, they search for periods constantly below a range of exogenous thresholds. Kapica et al. [38] compare energy droughts for solar PV and wind in Europe, either separately or as a combined portfolio, for historical and future weather conditions retrieved from eight regional climate models. Their identification method again draws on [19] with thresholds relative to seasonal and regional averages. Hu et al. [39] explore the risk of simultaneously occurring VRE droughts for European countries for both single technologies and a VRE portfolio, and compare it to droughts arising across all of Europe assuming perfect electricity exchange. Breyer et al. [40] model decarbonization pathways for Europe until 2040 or 2050, using the weather year 2005. They screen their model outcomes for drought periods constantly below a range of thresholds relative to the maximum availability of wind and PV separately and combined for Germany and Europe. Kaspar et al. [41] analyze wind and PV separately and combined for Germany. For given averaging intervals, they scan for moving averages below exogenously set thresholds. Mockert et al. [42] also study VRE droughts in Germany in the context of prevailing weather regimes and lean on Kaspar et al. [41] for drought identification. Mayer et al. [43] analyze concurrent PV and wind droughts in Hungary by identifying consecutive time steps with availability from both technologies below exogenous thresholds. Ohba et al. [44] link wind and solar droughts, both separately and combined as a portfolio, to four typical weather patterns arising in eastern Japan using an exogenously set threshold and neural network learning.
@@
In another strand of the literature, VRE shortage events are analyzed in the context of residual load. In addition to their VRE drought analysis, Raynaud et al. [19] investigate the imbalance of demand and VRE supply in Europe. Van der Wiel et al. [23] test whether low wind and solar generation as well as high residual events are linked to large-scale weather regimes arising over Europe. Otero et al. [49] investigate periods with simultaneously low wind and solar generation and high demand across 27 European countries based on absolute thresholds. Otera et al. [45] use copula analysis to analyze VRE droughts based on a threshold relative to a ten percent fraction of its mean availability, and residual load events above the 90% percentile of its distribution in Europe based on ERA5 reanalysis data. Allen & Otero [46] introduce standardized VRE droughts indices derived from meteorological indicators that incorporate both VRE generation and electricity demand. They showcase the proposed method for droughts in Europe. Ruhnau & Qvist [22] retrieve residual load time series from a model analysis of a least-cost decarbonized Germany power sector scenario and search for persistent positive residual load events. Dawkins & Rushby [8] characterize wind droughts and residual load events that qualify as droughts in the United Kingdom based on percentile thresholds. Francois et al. [50] explore positive residual load events (coined electricity deficits) for run-of-river hydropower and solar PV net of demand for Northern Italy, focusing on return periods lasting from single hours up to several years. Based on the assumption of sufficient storage availability, Plain et al. [51] identify low-solar periods using satellite data to size microgrids power systems in Africa based on the mismatch between solar production and demand. Bracken et al. [47] characterize VRE droughts across the U.S. concerning frequency, duration, magnitude, and seasonality given the existing energy infrastructure in 2020. They investigate single VRE technologies, VRE portfolios, and positive residual load events.
@@
Alternatively, the effect of extreme VRE conditions causing power system stress events can be detected based on electricity or energy system model outcomes. Grochowicz et al. [24] detect system stress events in Europe based on electricity shadow prices derived from a power sector model. They identify system-defining events that may last up to two weeks and account for a considerable share of total system costs. Such events occur between November and February and last between two and 13 days. Importantly, their findings are only applicable to the particular model setting and rely on a range of model-inherent parameter assumptions. Similar analyses have been carried out for different regions in the U.S., focusing on electricity prices and emissions [25], electricity prices and wind generation during winter storms [26] as well as electricity prices and system stability [48].
@@
While we do not aim to provide a complete review of the existing literature, we provide a structured overview of different approaches for identifying variable renewable energy droughts across multiple regions and technologies from the perspective of the energy system. Based on the literature review, we focus our analysis on threshold-based methods, including variable renewable energy droughts and positive residual load events. We discuss possible definitions of VRE droughts and identification methods, focusing on renewable availability time series. As variable renewable energy potentials vary across technologies, space, and time, we elaborate on how to enable comparability across these dimensions, which has not yet been considered in the literature. We further discuss how VRE drought analysis can be extended to also include electric load, i.e., how to evaluate extreme PRL. This allows for identifying compound low-VRE availability and high-demand events indicating the need for dispatchable capacity. Drawing on the proposed method for VRE availability time series, we discuss how to analyze PRL events and discuss essential points of criticism.
@@ 4.3 Methodological trade-offs and outlook
Such a trade-off also applies to the technological scope of VRE drought analysis. Aside from methodological assumptions inherent to the generation of variable renewable availability time series, e.g., such for reanalysis, insights from single-technology analysis are broadly applicable but less policy-relevant for future energy systems that rely on a mix of renewable technologies. Conversely, analyzing a VRE portfolio is more policy-oriented but involves energy modeling or exogenous assumptions on the composition of the investigated capacity mix. VRE capacity mixes determined by an energy model are optimal only in the context of the other optimized system components and flexibility options that are represented in the model (e.g., storage, electricity exchange, or direct-electrified applications from the heat, industry, or transport sector). Similarly, exogenous capacity mixes rely on user-specific assumptions. These aspects limit the validity of insights to investigated scenarios.
While we provide general guidance on the advantages and drawbacks of different VRE shortage identification methods, a thorough quantitative comparison would be desirable, ideally for a range of variable renewable technologies and electric load profiles for different world regions and weather years. In the Supplementary Information, we briefly indicate how a comprehensive comparative study could be carried out based on a stylized case study. Similar analyses have been conducted by Ohlendorf & Schill [35] and Potisomporn et al. [17] for a subset of methods and technologies as well as a limited number of regions. Another key aspect that requires attention is the choice of the threshold parameter, particularly in distinguishing between severe VRE droughts and prolonged periods of below-average availability. This distinction could support the selection of relevant weather years to enable weather- and climate-resilient energy system analyses and well-informed decarbonization policies. Additionally, assessing the sensitivity of outcomes to different VRE availability datasets would be useful and could be valuable feedback to the meteorological research community that provides and continuously improves such datasets [53].